A girl had designed a clap switch for a science exhibition that enabled her to switch on or off an alarm just
with clapping of hands. While testing her device in a hall she noticed that once the alarm has sounded it
followed with another one due to echo of the clap that the sound reflected by the walls. She recorded the two
soundings of alarm with her tape-recorder and found out that the time difference in between them is 0.10
seconds. If the distance of the walls be 16 m, what is the maximum possible speed of the sound

Grade:12th Pass

1 Answers

Sudheesh Singanamalla
13 years ago

Time of the sound to go hit the wall and then return back is 0.10 seconds.

so time taken by sound to just hit the wall is 0.10/2 = 0.05 seconds.

distance between the walls = 16m

 

we use the formula : speed = distance / time

which gives us the speed of the sound in air

speed = 16 / 0.05 = 1600/5 = 320 metres per second.

 

so the speed of sound in air is 320 metres per second.

 

Please approve !